# MobileDiffusion: Subsecond Text-to-Image Generation on Mobile Devices

### v1

@@ 3.2.2 Convolution Blocks
Separable convolution. Residual blocks in vanilla convolutional layers typically involve a substantial number of parameters, prompting various endeavors to enhance the parameter efficiency of convolutional layers. One proven approach in this context is separable convolution [16]. We have observed that replacing vanilla convolution layers with lightweight separable convolution layers in the deeper segments of the UNet yields similar performance. As a result, we replace all the convolutional layers in the UNet with separable convolutionas, except for the outermost level. The separable convolutional block we adopted, as illustrated in Figure 2, shares similarities with ConvNeXt [26] but employs a smaller $3\times 3$ kernel size. While we experimented with larger kernel sizes, such as $7\times 7$ and $9\times 9$, we found that they did not provide additional improvements.
This substitution enables us to reduce the model’s parameter count by approximately 10%.
@@
Prune redundant residual blocks. Convolution operations on high-resolution feature maps are especially computationally expensive, and pruning the convolution blocks is a straightforward way to improve model efficiency. Through comprehensive studies, we can effectively reduce the number of required residual blocks from the previous 22 (in SD) to a more efficient and streamlined 12. More specifically, we set 1 layer per block instead of 2 in SD, except for the innermost blocks. While maintaining a balance between resource consumption and model performance,
This pruning strategy optimizes computational efficiency by 19% and leads to 15% parameter reduction.
Models #Channels #ConvBlocks #(SA+CA) ##\##Params(M) ##\##GFLOPs Latency(ms) Model Size (GB)
SD-XL [36] (320, 640, 1280) 17 31+31 2,300 710 29.5 5.66
SD-1.4/1.5 (320, 640, 1280, 1280) 22 16+16 862 392 21.7 2.07
SnapFusion [23] (320, 640, 1280, 1280) 18 14+14 848 285 15.0 1.97
MobileDiffusion (320, 640, 1024) 11 15+18 386 182 9.9 1.04
MobileDiffusion-Lite (320, 640, 896) 11 12+15 278 153 8.8 0.82
Table 1: Comparison with other recognized latent diffusion models. Latency and GFLOPs, computed with $\mathsf{jit}$ per forward step, are measured for an input latent size of $64\times 64$ on TPU v3. Model size (fp16) includes all, i.e., UNet, text encoder and VAE decoder.
3.3 Architecture Candidates
@@
Employing a the potential model architecture optimization techniques discussed earlier, we meticulously refine our selection of viable architecture candidates by imposing upper bounds on the number of parameters and floating-point operations per second (FLOPs). Specifically, for our model, denoted as MobileDiffusion (MD), we set a limit of 400 million parameters and 200 GFLOPs to ensure computational efficiency without compromising model performance. Notably, for ease of model search, we combine the innermost level of the down stack, the up stack, and the middle bottleneck into a unified module, as they share the same feature map dimension. This corresponds to the 5 layers in the middle of MD architecture as illustrated in Figure 2. Within this constrained search space, we systematically explore variations in the number of transformer layers and innermost channel dimensions, concentrating on the pivotal architectural parameters that significantly influence model performance. Due to the computational demands associated with training diffusion models, we prioritize optimizing these critical parameters to enhance the model’s efficiency and effectiveness. In Table 1, we present a comparative analysis of our selected MD candidates, encompassing both the base model and a lightweight variant, against SD UNets and an on-device optimized SnapFusion UNet. The results underscore that our MD models exhibit a reduction in residual blocks, balanced attention layers, and enhanced parameter and runtime efficiency. Notably, the use of three blocks demonstrates superior visual attractiveness while maintaining comparable parameter counts, aligning with our earlier assertion that higher-level cross attention is integral to achieving high-quality diffusion models.
3.4 Sample Efficiency
In pursuit of sample efficiency, we employ off-the-self progressive distillation techniques [43, 32, 23]. The distillation method recursively condense the sampling trajectory of the teacher model into half number sampling steps. Our findings indicate that through distillation, a minimum of 8 steps can be achieved without compromising image quality.
We additionally adopt the recent advances of diffusion-GAN hybrid models, such as  [54, 56, 57]. In particular, we implement UFOGen [57] on our model. [57] proposes to finetune the pre-trained diffusion model with an hybrid objective of reconstruction term and adversarial term. This approach have demonstrated remarkable capabilities to transform a pre-trained diffusion model into a single-step generative model, pushing the sampling efficiency of diffusion models to a new level.
@@ 4 Experiments
SD 49.5 26.7 0.76 0.037 200
Original 39.3 30.0 0.83 0.032 125
Ours 9.8 30.2 0.84 0.032 65
Table 2: Decoder variants before and after finetuning.
We begin by providing an overview of the training specifics in Section 4.1, which encompass experimental setup, dataset, and the evaluation protocol. In Section 4.2, following that, we present the primary results of our text-to-image generation including both cfg-aware distilled and UFO finetuned models, including quantitative metrics, generated samples, perceptual assessments, as well as on-device benchmarks. Building upon the foundation of the pre-trained MobileDiffusion models, in Section 4.3, we delve into an exploration of various mobile applications, such as the integration of controls through plugins and of personalization capabilities via LoRA finetuning.
Table 3: Qualitative comparison between SD-XL, MD-Lite, and MD in DDIM 50 steps, as well as MD-distilled and MD-UFOGen.
4.1 Training
UNet We offer two versions of UNet as outlined in Table 1. When compared to SD-1.5, MobileDiffusion achieves a compression rate of 46.4%. Additionally, MobileDiffusion-Lite, characterized by a narrower model width and shallower transformer blocks, achieves even higher compression, reaching 29.9%. Both variants also demonstrate significant advantages in terms of inference efficiency.
Refer to caption Refer to caption
Input SD
Refer to caption Refer to caption
Original Ours
Figure 3: VAE reconstruction comparison among SD decoder, our original decoder and retrained efficient decoder.
VAE In our model, we use the VAE with some modifications over the SD’s VAE. Same as SD’s VAE, our VAE encoder downsamples the spatial resolution of the image by 8×8$\times$8 ×. Our encoder encodes a 3-channel RGB image into an 8-channel latent variable, instead of the 4 channels used in SD. This modification enhances the quality of image reconstruction, as demonstrated in Figure 3. Similar to other VAEs used for latent diffusion models, our VAE is trained with a combination of loss functions, including $\mathcal{L}_{2}$ reconstruction loss, KL divergence for regularization, perceptual loss, and adversarial loss. We train the VAE with batch size 256 for 2 million iterations using the same dataset for diffusion model training.
To further enhance efficiency, we design a lightweight decoder architecture by pruning the original one’s width and depth. We trained the lightweight decoder for 400K iterations with the encoder frozen. Since the encoder is frozen, we remove the unnecessary KL regularization in the objective. We also reduce the weight of adversarial loss. The resulting lightweight decoder leads to a significant performance boost, with nearly 50% latency improvement when decoding a $512\times 512$ image. Quantitative comparison on COCO2017 validation in Table 2 shows the lightweight decoder’s quality is on par with the original one.
Text Encoder We use OpenAI CLIP-ViT/L14 [37] as used in SD for text encoding, which is known to be powerful with a relatively compact size.
Dataset Our training process leverages a proprietary dataset, comprising an extensive collection of 150 million image-text pairs from public web [5]. The majority of images have a resolution greater than $256\times 256$, and 40 million images with a resolution over $512\times 512$. To prepare these images for training, we follow a consistent preprocessing methodology. Specifically, we crop the largest central square from each image and resize it to match the desired resolution.
Optimization Our models are trained with the AdamW optimizer [27], configured with a learning rate ($lr$) of 0.001, $\beta_{1}$ of 0.9, $\beta_{2}$ of 0.999, and a weight decay of 0.01. For image resolutions of $256\times 256$, we employed a batch size of 4,096, while for resolutions of $512\times 512$, the batch size was set at 2,048. The training process encompassed two phases: first, we trained at $256\times 256$ resolution for 0.75M steps to capture high-level semantics, followed by an additional 0.25M steps at a resolution of $512\times 512$ to refine features.
Training Cost When searching and selecting superior candidates, we have found that it suffices to rely on the FID and CLIP scores reported at the 30K iteration mark before which better designs can already produce text-image aligned results. This corresponds to approximately 4 hours of computation using 32 TPUs with 16GB memory each. We have designed a stopping mechanism for our experiments, which halts when the FID and CLIP scores show signs of slower improvement compared to previous candidates. This strategy works reasonable well in the process. In total, our endeavor consumes approximately 512 TPUs spanning 30 days to complete the refinement of MDs. The training duration for both MD and MD-Lite spans approximately 5 days in total.
4.2 Main Results
We present comprehensive results of the efficiency and quality of text to image generation, spanning quantitative and qualitative comparisons, and on-device benchmarks.
Quantitative Evaluation In Table 4, we conducted a comprehensive comparison of various text-to-image generative models. We prioritized sample efficiency by employing both cfg-aware distillation [23] and UFOGen [57] fine-tuning techniques, and we present the corresponding numerical results. For DDIM and the distilled models, we achieved the lowest FID scores by adjusting the cfg scales around 3.0. Since UFOGen does not involve cfg, we report the FID scores as is. We replicated the training settings of SD to match those of MD for a fair comparison. Our primary focus is on comparing our models using various samplers and the SD-replicated baseline, as FID evaluation can be subject to noise variations across different research works. The MD DDIM achieves a performance level close to that of the baseline SD-replicated model while being 50% smaller and 50% faster. Comparing the distilled 8-step model and the UFOGen finetuned 1-step model, we observe that their FID scores are comparable. Additionally, we compute the CLIP-ViT-g/14 score, and it’s worth noting that both models can achieve around 0.33. This further highlights the effectiveness of our approach in achieving sample efficiency.
@@ 4.2 Main Results
GigaGAN [18] GAN 1 9.09 0.9 0.98
LAFITE [62] GAN 1 26.94 0.23 0.003
Parti [59] AR - 7.23 20.0 5.00
[…]
GLIDE [34] DDPM 250 12.24 5.00 0.25
Imagen [42] DDPM 256 7.27 3.60 0.45
SD [39] DDIM 50 8.59 0.86 0.60
SnapFusion [23] Distilled 8 13.5 0.85 -
PIXART-$\alpha$ [4] DPM 20 10.65 0.6 0.025
BK-SDM [21] DDIM 50 16.54 0.50 -
SD-replicated [^1] We replicated SD on our training data and pipeline. DDIM 50 8.43 0.86 0.15
MD DDIM 50 8.65 0.40 0.15
Distilled 8 9.01
UFOGen 1 11.67
MD-Lite DDIM 50 9.45 0.26 0.15
Distilled 8 9.87
UFOGen 1 12.89
Table 4: Quantitative evaluations on zero-shot MS-COCO.
Qualitative Comparison In Table 3, we present a visual comparison between SD-XL, MD, distilled MD with 8 steps, and UFOGen finetuned with 1 step. Notably, MD consistently produces high-quality results across various types of samplers, even when compared to the strong baseline SD-XL. This demonstrates the potential for on-device applications. More examples can be found in Appendix.
On-device Benchmark We conducted benchmarking of our proposed MD using tools [^2] https://github.com/apple/ml-stable-diffusion/tree/main. on iPhone 15 Pro. As depicted in Table 5, MD exhibits superior efficiency in various aspects, including the text encoder, VAE decoder, UNet per-step inference, and the resulting overall latency, for both the distilled 8-step and finetuned 1-step models.
@@
SnapFusion [23] [^3] Numbers are from [23], measured on iPhone 14 Pro. 4 116 230 8 1960
UFOGen 4 285 1580 1 1869
MD 4 92 142 8 1232
1 238
MD-Lite 4 92 123 1 219

### v2

@@ 3.3 Optimize convolution
Separable convolution. Residual blocks in vanilla convolutional layers typically involve a substantial number of parameters, prompting various endeavors to enhance the parameter efficiency of convolutional layers. One proven approach in this context is separable convolution [19]. We have observed that replacing vanilla convolution layers with lightweight separable convolution layers in the deeper segments of the UNet yields similar performance. As a result, we replace all the convolutional layers in the UNet with separable convolutions, except for the outermost level. The separable convolutional block we adopted, as illustrated in Figure 2, shares similarities with ConvNeXt [31] but employs a smaller $3\times 3$ kernel size. While we experimented with larger kernel sizes, such as $7\times 7$ and $9\times 9$, we found that they did not provide additional improvements.
@@
Prune redundant residual blocks. Convolution operations on high-resolution feature maps are especially computationally expensive, and pruning is a straightforward way to improve model efficiency. Through a comprehensive network search, we reduce the number of required residual blocks from 22 (in SD) to more efficient and streamlined 12. More specifically, we set 1 layer per block instead of 2 in SD, except for the innermost blocks. While maintaining a balance between resource consumption and model performance.
Models #Channels #ConvBlocks #(SA+CA) ##\##Params(M) ##\##GFLOPs
SD-XL [42] (320, 640, 1280) 17 31+31 2,600 710
SD-1.4/1.5 [45] (320, 640, 1280, 1280) 22 16+16 862 392
SnapFusion [27] (320, 640, 1280, 1280) 18 14+14 848 285
DiT XL/2 [41] (1152) 0 28+0 675 525
MobileDiffusion (320, 640, 1024) 11 15+18 386 182
Table 1: Comparison with other recognized latent diffusion models.
3.4 Model details
@@
Employing the optimization techniques discussed earlier, we meticulously refine our selection of viable architecture candidates by imposing upper bounds on the numbers of parameters and FLOPs. We aim at 400 million parameters and 200 GFLOPs to achieve instant generation on devices. Notably, for the ease of model search, we combine the innermost level of the down stack, the up stack, and the middle bottleneck into a unified module, as they share the same feature map dimension. This corresponds to the 5 layers in the middle of architecture as illustrated in Figure 2. Within this constrained search space, we systematically explore variations in the number of transformer layers and innermost channel dimensions, concentrating on the pivotal architectural parameters that significantly influence model performance. Due to the computational demands associated with training diffusion models, we prioritize optimizing these critical parameters to enhance the model’s efficiency and effectiveness. In Table 1, we present a comparative analysis of MobileDiffusion, against SD UNets, SnapFusion UNet [27] optimized for on-device, and DiT XL/2 [41]. It underscores that MobileDiffusion exhibits the highest efficiency. MD conveys more computation from convolution to attention than UNets, and embodies fewer FLOPs than pure transformer architectures.
Refer to caption Refer to caption Refer to caption Refer to caption
Input SD Ours Ours-distilled
Figure 3: VAE reconstruction comparison among SD decoder, our decoder and our distilled decoder.
3.5 Optimize VAE
In latent diffusion, VAE converts a $H\times W\times 3$ image to a $\frac{H}{f}\times\frac{W}{f}\times c$ latent. SD [45] uses $f=8$ and $c=4$, while EMU [9] uses $f=8$ and $c=16$. Using smaller channel number $c$ results in better compression but worse reconstruction. After an empirical study, we choose $f=8$ and $c=8$ for training our VAE. This modification enhances the quality of image reconstruction, as demonstrated in Figure 3. Similar to other VAEs used for latent diffusion models, our VAE is trained with a combination of loss functions, including $\mathcal{L}_{2}$ reconstruction loss, KL divergence for regularization, perceptual loss, and adversarial loss. We train the VAE with batch size 256 for 2 million iterations using the same dataset for diffusion model training. As pointed out in [27], when evaluating with a smaller number of diffusion steps, image decoder becomes a non-negligible component. To further enhance efficiency, we design a lightweight decoder architecture by pruning model width and depth. We train the lightweight decoder for 400K iterations with the encoder frozen. We remove the unnecessary KL regularization in the objective, and reduce the weight of adversarial loss. The distilled decoder leads to a significant performance boost, 3 times faster than SD’s decoder as shown in Table 4.
4 Elucidating the design space of UFOGen
In this section, we detail our studies on the ultra-fast one-step diffusion models, focusing on enhancing image quality and, crucially, enabling the integration of a broad spectrum of downstream applications into the one-step generation paradigm. Our investigation primarily concentrates on UFOGen [62], an innovative adversarial fine-tuning approach that has shown considerable promise for one-step text-to-image generation. We start with a concise overview of UFOGen, followed by a discussion on the series of ablation studies we undertook to refine its training methodology
4.1 Overview of UFOGen
UFOGen employs adversarial fine-tuning of a pre-existing diffusion model to achieve one-step generation. This process involves the one-step generator $G_{\theta}$, being trained jointly with a discriminator $D_{\phi}$ with the following objective
$\min_{\theta}\max_{\phi}\mathbb{E}_{q(x_{0})q(x_{t-1}|x_{0}),p_{\theta}(x_{0}^{\prime})p_{\theta}(x_{t-1}^{\prime}|x_{0}^{\prime})}\Bigl{[}$ (1)
$\underbrace{[\log(D_{\phi}(x_{t-1},t))]+[\log(1-D_{\phi}(x^{\prime}_{t-1},t))]}_{\text{adversarial loss}}+\underbrace{\lambda\gamma_{t}\left\lVert x_{0}-x_{0}^{\prime}\right\rVert^{2}}_{\text{diffusion loss}}\Bigl{]},$
where $q(x_{0})$ is the training data distribution, $q(x_{t-1}|x_{0})$ is the forward diffusion process. The generator is parameterized by generating a clean image $x_{0}^{\prime}=G_{\theta}(x_{t},t)\sim p_{\theta}(x_{0}^{\prime})$, and $x_{t-1}^{\prime}\sim p_{\theta}(x_{t-1}^{\prime}|x_{0}^{\prime})$ is applying forward diffusion process on generator’s output $x_{0}^{\prime}$. Besides the adversarial loss, UFOGen also includes the original diffusion loss in its objective, and this is shown to be crucial for stabilizing the training. The generator and discriminator both have the same structure and initialized by the pre-trained diffusion model. The UNet discriminator aggregates logits across output locations make decisions.
UFOGen has demonstrated its efficacy in generating high-quality images from text prompts in a single step, significantly enhancing the efficiency of sampling in diffusion models. However, training strategies of UFOGen have not been well-studied. In addition, seamlessly incorporating the separately-trained modules such as LoRA [20], ControlNet [65] and T2I-adapter [39] with one-step sampling remains unexplored. While the original UFOGen study [62] showcases successful image generation from depth maps and canny edges, these outcomes are achieved through task-specific adversarial training, necessitating a separate copy for each new conditional task — far from the ideal scenario. To this end, we carefully studied the design space of UFOGen, drawing motivations from related methods. Our finalized design choice is obtained by comparing different variants on both the text-to-image generation and downstream conditional generation tasks.
4.2 Reconstruction term in the training objective
The first design aspect worth exploring is the role of the reconstruction term in UFOGen’s training objective. The formulation of UFOGen, as specified in Equation 1, integrates both adversarial and reconstruction losses. The purpose of the adversarial loss is to enhance the visual accuracy of images produced through a singular diffusion step. Traditionally, a diffusion model’s single-step generation process is calibrated to forecast the expected pristine image from a given noisy input, symbolized by $\mathbb{E}[x_{0}|x_{t}]$. As a result, this often yields images that are somewhat blurry, lacking in fine details due to the inherent averaging across the data distribution. The inclusion of an adversarial loss term empowers the model to generate sharper images, particularly at elevated noise levels. Conversely, the functionality of the reconstruction loss within the training regime is more nuanced. It is postulated to act as a stabilizing force during the adversarial training phase, offering a straightforward learning directive. Beyond this, we propose that it plays a vital role in maintaining the integrity of the pre-trained diffusion model’s characteristics. Such preservation is crucial; in its absence, the one-step model risks deviating markedly from the original feature space of the pre-trained model. Such a divergence could severely limit the compatibility with downstream modules specifically tailored for the original diffusion framework.
Our conjecture suggests it might be helpful to provide enhanced regularization to encourage the preservation the pre-trained model’s characteristics, ensuring that the one-step model remains closely aligned with the foundational diffusion model’s feature landscape. To this end, we examined some alternative options to the reconstruction loss in Equation 1. The first one is a distillation loss that explicitly align the features of UFOGen’s generator to be aligned with the teacher model:
$\mathcal{L}_{\text{distill}}=\left\lVert\text{sg}\left(G^{\text{teacher}}(x_{t},t)\right)-x_{0}^{\prime}\right\rVert^{2},$ (2)
where $G^{\text{teacher}}_{\theta}$ is the pre-trained teacher model, and $\text{sg}(\cdot)$ stands for stop gradient. Note that the distill loss is similar to the regularization term in the objective of Adversarial Diffusion Distillation (ADD) [49].
More generally, we also test the EMA distillation objective, which is formalized as follows
$\mathcal{L}_{\text{ema}}=\left\lVert\text{sg}\left(G^{\text{EMA}}_{\theta}(x_{t},t)\right)-x_{0}^{\prime}\right\rVert^{2},$ (3)
where $G^{\text{EMA}}_{\theta}$ is the exponential moving average of $G_{\theta}$. Given that $G_{\theta}$ originates from a pre-trained model, the EMA mechanism effectively retains a substantial portion of the pre-trained model’s information. This loss can also ensure that the essence of the pre-trained model is conserved, while providing more flexibility. Note that the distillation loss in Equation 2 is a special case of our loss, with the EMA decay parameter being set to 1. We conduct comprehensive ablation studies on these loss terms in Section 5.3.
4.3 Parameter-efficient adversarial fine-tuning
Current methods in adversarial diffusion fine-tuning, exemplified by UFOGen [62] and ADD [49], adjust the entire model. This extensive parameter modification can lead to significant shifts in the internal feature representation, potentially undermining the model’s ability to work effectively with downstream tasks. Parameter-efficient fine-tuning emerges as a compelling solution, drawing from its successful application across both the visual and language fields [20, 11]. Among these methods, LoRA (Low-Rank Adaptation) fine-tuning [20] stands out for its ability to refine diffusion models [36], offering a balanced approach that supports rapid sampling while maintaining task adaptability.
Motivated by the success of LCM-LoRA, we explore the potential of integrating LoRA fine-tuning within the UFOGen training framework. We apply LoRA exclusively to the generator, initializing the model weights from the pre-trained diffusion framework while setting the LoRA layers to initialize with a null effect. The discriminator is still initialized from the pre-trained diffusion model.
@@ 5 Experiments
We begin by providing an overview of the training specifics in Section 5.1, which encompass experimental setup, dataset, and the evaluation protocol. In Section 5.2, we present the primary results of our text-to-image generation including quantitative metrics, generated samples, perceptual assessments, as well as on-device benchmarks. In Section 5.4, we delve into an exploration of various mobile applications, such as the integration of controls through plugins and of personalization capabilities via LoRA finetuning.
5.1 Training Details
Dataset Our training process leverages a proprietary dataset, comprising an extensive collection of 150 million image-text pairs from public web [6]. The majority of images have a resolution greater than $256\times 256$, and 40 million images with a resolution over $512\times 512$. To prepare these images for training, we follow a consistent preprocessing methodology.
Optimization Our models are trained with the AdamW optimizer [32], configured with a learning rate of 0.001, $\beta_{1}$ of 0.9, $\beta_{2}$ of 0.999, and a weight decay of 0.01. We adopt a progressive training strategy: first, we train at $256\times 256$ resolution with batch size 4096 for 0.75M steps to capture high-level semantics, followed by an additional 0.25M steps at a resolution of $512\times 512$ with batch size 2048 to refine features.
Training Cost When searching and selecting superior candidates, we rely on the FID and CLIP scores reported at 30K training steps to decide the candidates to proceed. This corresponds to approximately 4 hours of computation using 32 TPUs with 16GB memory. We have designed a stopping mechanism for our experiments, which halts when the FID and CLIP scores show signs of slower improvement compared to previous candidates. This strategy works reasonable well in the process. In total, our endeavor consumes approximately 512 TPUs spanning 15 days to complete the network search.
5.2 Text-to-Image Generation
We present comprehensive results of the efficiency and quality of text-to-image generation, spanning quantitative and qualitative comparisons, and on-device benchmarks.
Quantitative Evaluation In Table 2, we conducted a comprehensive comparison of various text-to-image generation models. For DDIM, we achieved the lowest FID scores by adjusting the cfg scales around 3.0. Since UFOGen does not have cfg, we report the FID scores as is. We also compare CLIP scores of models close to ours. While being compact in model size, our MobileDiffusion achieves better metrics than previous solutions of SnapFusion [27] and UFOGen [62]. The figure on the right shows the FID vs. CLIP curve by comparing MobileDiffusion and Stable Diffusion v1.5 using 50-step DDIM by varying cfg.
We also consider human preference evaluation using the standard HPS v2 benchmark [58]. We compare with SD v2.0 and v1.4 in Table 3. Despite with much smaller model size and number of inference steps, our model obtained comparable performance with Stable Diffusion variants.
@@ 5.2 Text-to-Image Generation
GigaGAN [22] 1-step 1 9.09 - 0.9 0.98
LAFITE [66] 1-step 1 26.94 - 0.23 0.003
@@
Imagen [47] DDPM 256 7.27 - 3.60 0.45
SD [45] DDIM 50 9.62 0.304 0.86 0.60
PIXART-$\alpha$ [5] DPM 20 10.65 - 0.6 0.025
BK-SDM [24] DDIM 50 16.54 - 0.50 -
SnapFusion [27] Distilled 8 13.5 0.308 0.85 -
UFOGen [62] 1-step 1 12.78 0.317 0.86 0.6
MoibleDiffusion DDIM 50 8.65 0.325 0.39 0.15
1-step 1 11.67 0.320
Table 2: Quantitative evaluations on zero-shot MS-COCO 2014 validation set (30K).
Models Anim. Concept Painting Photo Avg
SD-v2.0 (50 steps) 27.48 26.89 26.86 27.46 27.17
SD-v1.4 (50 steps) 27.26 26.61 26.66 27.27 26.95
MD (50 steps) 27.52 27.13 27.30 27.26 27.30
MD (1 step) 27.05 26.33 26.41 26.80 26.65
Table 3: HPS v2 benchmark.
On-device Benchmark We conducted benchmarking of our proposed MD using tools [^1] https://github.com/apple/ml-stable-diffusion/tree/main. on iPhone 15 Pro. As depicted in Table 4, MD exhibits superior efficiency in various aspects, including the text encoder, VAE decoder, UNet per-step inference, and the resulting overall latency.
@@
SD 1.5 [45] 4 285 357 20 7429
SnapFusion [27] 4 112 203 8 1740
UFOGen [62] 4 285 357 1 646
MD-UFO 4 92 142 1 238
@@
5.3 Ablation on UFOGen design choices
In this section, we ablate the design choices of UFOGen discussed in Section 4. In particular, we tried different formulations for reconstruction loss and compared LoRA fine-tuning versus full parameter fine-tuning. We report the FID scores in Table 5. Our ablation results indicate that
  • LoRA adversarial fine-tuning consistently performs worse than full parameter fine-tuning
  • All reconstruction losses are effective in stabilizing the adversarial fine-tuning. The distill loss obtains slightly better performance.
Reconstruction loss No LoRA LoRA
Diffusion loss 11.67 13.45
Distill loss 11.20 13.24
EMA distill loss 12.08 14.05
Table 5: Ablation on different adversarial finetuning settings.
Despite the slight advantage of using distill loss reflected in the FID metric, we conduct more in-depth study on the impact of training design choices on downstream application. Due to the constraint of space, we provide discussion on the study in the Appendix. We observe that, surprisingly, the original diffusion loss has the best results on applications. Therefore, our comprehensive exploration of the design space of UFOGen’s training suggest that the original setting, namely full parameter fine-tuning with diffusion loss, is the best recipe for one-step diffusion model. As a result, we adopt this setting throughout the paper.